# Mechanically concealed holes

### v1

@@ Abstract
When a hole is introduced into an elastic material, it will usually act to reduce the overall mechanical stiffness. A general ambition is to investigate whether a stiff shell around the hole can act to maintain the overall mechanical properties. We consider the basic example situation of an isotropic, homogeneous, linearly elastic material loaded uniformly under plane strain for low concentrations of holes. As we demonstrate, the thickness of the shell can be adjusted in a way to maintain the overall stiffness of the system. We derive a corresponding mathematical expression for the thickness of the shell that conceals the hole. Thus, one can work with given materials to mask the presence of the holes. One does not necessarily need to adjust the material parameters and thus materials themselves. Our predictions from linear elasticity continuum theory are extended to atomistic level using molecular dynamics simulations of a model Lennard-Jones solid. Small deviations from linear elasticity theory can be minimized by tuning the hole-to-system size ratio in the molecular dynamics simulations. This extension attests the robustness of our continuum predictions even at atomistic scales. The basic concept is important in the context of light-weight construction.
@@
Introduction. Saving resources and fuel is a major concern of recent production lines and construction design. Besides aspects of sustainability, pure economic reasons favor corresponding achievements. Saving materials and energy reduces overall costs [1]. Therefore, light-weight construction remains key to future technological developments. Nature provides corresponding examples, maybe bones being the most obvious ones [2, 3]. Thanks to their stiff structure they provide overall stability for the whole organism, yet the many cavities of various types of bones reduce their overall weight. This structure saves energy of motion and increases agility and mobility.
@@
Analogously, for many components of machines, vehicles, aircrafts, or other devices, reducing weight provides significant benefit [4, 5]. However, the overall design has often been developed for years or decades and been adjusted to near perfection. In such cases, changing the dimension or shape of individual components to reduce their weight, or others of their mechanical properties like stiffness, provides additional challenges.
@@
Therefore, we focus on the idea of introducing holes into materials to save weight [6, 7]. Advanced strategies of designing materials with cylindrical or spherical holes had been already found beneficial for this purpose using approaches of topological optimization [8]. In our case, holes are considered to be mechanically masked in a way so that their presence is not noted on the overall, macroscopic scale of the material. Still, the overall mechanical stiffness shall be maintained. Together, this concept results in a component of identical mechanical properties, yet of reduced weight. Key is to introduce holes (cavities) that are surrounded by stiffer shells so that the overall, combined mechanical stiffness is the same as in the absence of the holes.
@@
The model geometry of introducing hollow cylindrical shells in three-dimensional solids (or hollow circular rings in two-dimensional solids) has also been applied to study diverse other problems, ranging from elastostatics [9, 10] to cavitation in soft solids [11, 12]. Amongst these, the idea of mechanically masking holes inside a solid, in fact, has been studied before in terms of “mechanical cloaking”. It provides the solid with a property of “mechanical unfeelability” of the holes within, in terms of the overall macroscopic response. Additionally, mechanically masking holes and corresponding optimization has been realized to conceal mechanical [13], thermo-mechanical [14], as well as dispersive properties [15] of solids, especially in the context of metamaterials design [16, 17, 18]. Linear elastodynamic cloaking of cylindrical holes was discussed for infinitesimal in-plane deformations [19].
@@
Recently, mechanical concealment in elastostatic situations received additional attention. Notably, when viewed from the perspective of optimized materials design, elastostatic cloaking has been outlined following two strategies. Either the displacement fields [20] or the elastic moduli [21] and thus the type of employed materials were used as design parameters. Interestingly, a recent analytical and numerical work [22] showed that elastostatic mechanical cloaking of a circular inclusion in two-dimensional geometries can be attained by coating the inclusion using several concentrically arranged circular rings and tuning their shear moduli. This strategy eliminates the complexity of additionally tuning the Poisson ratio.
@@
To this end, we present an alternative strategy to illustrate the concept of mechanically concealed holes using a basic example situation. We address uniform loading under plane-strain conditions of a material that is homogeneous, isotropic, linearly elastic, and, in principle, infinitely extended. The cylindrical holes that we introduce are of sufficiently low concentration so that we may neglect their mutual mechanical interaction. In reality, we may, for instance, think of cylindrical holes drilled into plates or blocks of material. Using analytical calculations as well as molecular dynamics simulations, we demonstrate that, indeed, under such conditions, we can introduce stiff cylindrical shells around the cylindrical holes so that the presence of the holes is effectively masked and mechanically concealed. Our sole tuning parameter is the thickness of the shell surrounding the hole.
Theoretical background. We start our theoretical consideration from the basic theory of linear elasticity [23]. Stress $\bm{\sigma}^{\mathrm{(3d)}}$ and strain $\bm{\varepsilon}^{\mathrm{(3d)}}$ are related to each other via the shear modulus $\mu$ and the Lamé parameter $\lambda$, which is associated with compressibility,
@@
Derivation of the mechanical solution for a shelled hole. We address the situation in polar coordinates. Our coordinate system is centered in the hole, see Fig. 1. The hole has a radius $a$, while the surrounding shell of inner radius $a$ is of outer radius $b>a$. While the shell is of shear modulus $\mu^{\mathrm{i}}$ and Poisson ratio $\nu^{\mathrm{i}}$, the corresponding parameters of the surrounding elastic material are $\mu^{\mathrm{o}}$ and $\nu^{\mathrm{o}}$. Here and in the following, superscript “i” indicates the “inner” elastic material (shell), while “o” denotes the “outer” elastic material.
@@
Since we consider uniform loading, the stress tensor at infinite distance from the hole can be denoted as $\bm{\sigma}(r\rightarrow\infty)=-P\mathbf{I}$, where $P>0$. In this circularly symmetric situation, for linearly elastic systems, there is no angular dependence of the results on the polar angle $\varphi$. Thus, from all possible terms contributing to the stress function $F$ that satisfy Eq. (7) [24], we retain only those that do not imply any dependence on $\varphi$ in the physical solution. We formulate the stress function separately for the inner and outer regions,
@@
Thickness of the shell for mechanical concealment. We now turn to the central point. The shell around the hole shall mechanically conceal and mask the hole in a way that its presence is not noted from outside. This concealment shall be realized for given materials, that is, we may not modify the material parameters $\mu^{\{\mathrm{i,o}\}}$ and $\nu^{\{\mathrm{i,o}\}}$. For that purpose, we must choose the outer radius $b$ of the shell in a way that the mechanical solution outside the shelled hole appears in the same way as if the hole were not present at all.
@@
If the Poisson ratio of the shell is smaller compared to the surrounding material, $\nu^{\mathrm{i}}<\nu^{\mathrm{o}}$, the necessary outer radius and thus thickness of the shell diverges even with shells stiffer than the outside elastic material, $\mu^{\mathrm{i}}>\mu^{\mathrm{o}}$, if the shells are not overly stiff, see the red curves in Fig. 2. In such cases, for given materials, mechanical shielding will be challenging, and a stiffer choice of shell material may still be advisable. For example, the combination $\nu^{\mathrm{i}}=0.3$ and $\nu^{\mathrm{o}}=0.49$, see Fig. 2(a), depicts a compressible shell such as steel [25, 26] within a nearly incompressible background like a rubbery elastomer [25, 26]. However, given that steels are generally several orders of magnitude stiffer than elastomers, this does not represent an actual constraint for mechanical concealment. It would still be possible already using very thin shells.
@@
Thus, where possible, selecting most beneficial materials provides advantages. Leveraging the existence of various unconventional materials, one may even opt for materials of near-zero or negative Poisson ratios [27, 28, 29, 30, 31, 32, 33] to achieve advanced mechanical shielding. However, once a choice is made or the materials are determined by other factors, then still for many practical material combinations the strategy of simply adjusting the thickness of the shell provides an effective, realizable way for mechanical concealment.
@@
Molecular dynamics simulations. To proceed a step beyond continuum considerations, we seek to investigate whether our analytical solution for mechanical concealment of a hole carries over to atomistic level. For this purpose, we analyze a typical microscopic model solid using molecular dynamics (MD) simulations. We set combinations of $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$ and $b/a$ and test whether mechanical concealment can be achieved.
@@
In fact, atomistic investigations of mechanically concealed holes inside a solid are rare and challenging. Specifically, stabilizing solids with voids or cavities on the atomic scale often needs many-body terms in the interatomic interaction potentials. They come at a cost of losing simple, transparent models. Often, the descriptions invoke anisotropy due to angular dependence of the potentials. Our scope is to circumvent this issue and work with a description as simple as possible. Therefore, we use a two-dimensional, truncated and shifted Lennard-Jones (LJ) potential and adjust its length and energy scale, that is, the depth of the energy well, see the supplemental material for details [34], even if resulting energy wells are significantly deeper than for standard rare-gas solids. They imply very high mechanical moduli of the order of TPa [34]. As an advantage, we can work with a very simple and accessible potential and thus mimic a “toy model” for a strongly cohesive solid. As a further benefit, the LJ potential in two dimensions leads to stable hexagonal solids. The linearly elastic properties of solids of hexagonal symmetry are the same as for isotropic ones from a continuum perspective [23]. Therefore, our atomistic approach is in line with the continuum considerations above for isotropic elastic materials.

### v2

@@ Abstract
When a hole is introduced into an elastic material, it will usually act to reduce the overall mechanical stiffness. A general ambition is to investigate whether a stiff shell around the hole can act to maintain the overall mechanical properties. We consider this effect from a macroscopic continuum perspective down to atomistic scales. First, we focus on the basic continuum example situation of an isotropic, homogeneous, linearly elastic material loaded uniformly under compressive plane strain for low concentrations of holes. As we demonstrate, the thickness of the shell can be adjusted in a way to maintain the overall stiffness of the system. We derive a corresponding mathematical expression for the thickness of the shell that conceals the hole. Thus, one can work with given materials to mask the presence of the holes simply by adjusting the thickness of the surrounding shells, with no need to change the materials. Our predictions from linear elasticity continuum theory are extended to atomistic levels using molecular dynamics simulations of a model Lennard-Jones solid. These extensions attest the robustness of our predictions down to atomistic scales. Thus, they open a straightforward possibility to adjust the strategy of mechanical cloaking via atomistic manipulations. From both perspectives, the underlying concept is important in the context of light-weight construction.
I Introduction
@@
Saving resources and fuel is a major concern of recent production lines and construction design. Besides aspects of sustainability, pure economic reasons favor corresponding achievements. Saving materials and energy reduces overall costs König, Mathieu, and Vielhaber (2024). Therefore, light-weight construction remains key to future technological developments. Nature provides corresponding examples, maybe bones being the most obvious ones Lai et al. (2015); Yu et al. (2020). Thanks to their stiff structure they provide overall stability for the whole organism, yet the many cavities of various types of bones reduce their overall weight. This structure saves energy of motion and increases agility and mobility.
@@
Analogously, for many components of machines, vehicles, aircrafts, or other devices, reducing weight provides significant benefit Koffler and Rohde-Brandenburger (2010); Timmis et al. (2015). However, the overall design has often been developed for years or decades and been adjusted to near perfection. In such cases, changing the dimension or shape of individual components to reduce their weight, or others of their mechanical properties like stiffness, provides additional challenges.
@@
Therefore, we focus on the idea of introducing holes into materials to save weight Wang and Salhab (2009); Zhou et al. (2022). Advanced strategies of designing materials with cylindrical or spherical holes had been already found beneficial for this purpose using approaches of topological optimization Bruggi et al. (2022). In our case, holes are considered to be mechanically masked in a way so that their presence is not noted on the overall, macroscopic scale of the material. Still, the overall mechanical stiffness shall be maintained. Together, this concept results in a component of identical mechanical properties, yet of reduced weight. Key is to introduce holes (cavities) that are surrounded by stiffer shells so that the overall, combined mechanical stiffness is the same as in the absence of the holes.
@@
The model geometry of introducing hollow cylindrical shells in three-dimensional solids (or hollow circular rings in two-dimensional solids) has also been applied to study diverse other problems, ranging from elastostatics Lu, Xu, and Zhang (2017); Liu (2006) to cavitation in soft solids Kang, Wang, and Tan (2018); Kang and Tang (2021). Amongst these, the idea of mechanically masking holes inside a solid, in fact, has been studied before in terms of “mechanical cloaking”. It provides the solid with a property of “mechanical unfeelability" of the holes within, in terms of the overall macroscopic response. Additionally, mechanically masking holes and corresponding optimization has been realized to conceal mechanical Cheng et al. (2023), elasto-mechanical Bückmann et al. (2014); Diatta and Guenneau (2014); Achaoui et al. (2020); Stenger, Wilhelm, and Wegener (2012), thermo-mechanical Cheng et al. (2025), as well as dispersive properties Meirbekova and Brun (2020) of solids, especially in the context of metamaterials design Bückmann et al. (2015); Wang et al. (2022); Sanders, Aguiló, and Paulino (2021); Liu et al. (2025); Craster et al. (2023).
Polar materials had been employed as elastostatic cloaks Nassar, Chen, and Huang (2018); Zhang et al. (2024), often with functionally graded multilayered lattices Xu et al. (2020). Static mechanical cloaking of voids had also been realized recently via a design mechanism of generating irregular structures Yang et al. (2025). The paths to identify appropriate mechanical cloakings is frequently based on advanced transformation methods. They lead to complex structures of the metamaterials design and involve an elevated number of optimized subunits. For instance, the mechanical cloaking of a hexagonal bimode structure essentially requires stress optimization of the double-trapezoid units in the considered region Hai, Zhao, and Meng (2018).
@@
Although linear elastodynamic cloaking of cylindrical holes was discussed for infinitesimal in-plane deformations Yavari and Golgoon (2019), mechanical concealment in elastostatic situations still receives growing attention. Notably, when viewed from the perspective of optimized materials design, elastostatic cloaking has been outlined following two strategies. Either the displacement fields Fachinotti, Peralta, and Albanesi (2018) or the elastic moduli Sozio, Shojaei, and Yavari (2023) and thus the type of employed materials were used as design parameters. Interestingly, a recent analytical and numerical work Fielding (2024) showed that elastostatic mechanical cloaking of a circular inclusion in two-dimensional geometries can be attained by coating the inclusion using several concentrically arranged circular rings and tuning their shear moduli. This strategy eliminates the complexity of additionally tuning the Poisson ratio.
@@
To this end, we present an alternative strategy to illustrate the concept of mechanically concealed holes using a basic example situation, keeping the basic geometry as simple as possible. We start from continuum elasticity theory and address uniform static loading under plane-strain conditions of a material that is homogeneous, isotropic, linearly elastic, and, in principle, infinitely extended. The cylindrical holes that we introduce are of sufficiently low concentration so that we may neglect their mutual mechanical interaction. In reality, we may, for instance, think of cylindrical holes drilled into plates or blocks of material. Using analytical calculations, we demonstrate that, indeed, under such conditions, we can introduce stiff cylindrical shells around the cylindrical holes so that the presence of the holes is effectively masked and mechanically concealed. Our sole tuning parameter is the thickness of the shell surrounding the hole. The concept is then extended to atomistic scales using molecular dynamics simulations of an appropriate Lennard-Jones model solid. It maintains valid for discrete particulate structures. We mention that, in reality, colloids often provide example systems to visually bridge between atomistic and macroscopic scales and may serve for corresponding illustrations Herlach et al. (2010); Royall et al. (2024); Menath et al. (2023).
II Continuum elasticity solution
II.1 Theoretical background
We start our theoretical consideration from the basic theory of linear elasticity Landau and Lifshitz (1999). Stress $\bm{\sigma}^{\mathrm{(3d)}}$ and strain $\bm{\varepsilon}^{\mathrm{(3d)}}$ are related to each other via the shear modulus $\mu$ and the Lamé parameter $\lambda$, which is associated with compressibility,
@@ II.1 Theoretical background
II.2 Derivation of the mechanical solution for a shelled hole
We address the situation in polar coordinates. Our coordinate system is centered in the hole, see Fig. 1. The hole has a radius $a$, while the surrounding shell of inner radius $a$ is of outer radius $b>a$. While the shell is of shear modulus $\mu^{\mathrm{i}}$ and Poisson ratio $\nu^{\mathrm{i}}$, the corresponding parameters of the surrounding elastic material are $\mu^{\mathrm{o}}$ and $\nu^{\mathrm{o}}$. Here and in the following, superscript “i” indicates the “inner” elastic material (shell), while “o” denotes the “outer” elastic material.
@@ II.2 Derivation of the mechanical solution for a shelled hole
Since we consider uniform loading, the stress tensor at infinite distance from the hole can be denoted as $\bm{\sigma}(r\rightarrow\infty)=-P\mathbf{I}$, where $P>0$. In this circularly symmetric situation, for linearly elastic systems, there is no angular dependence of the results on the polar angle $\varphi$. Thus, from all possible terms contributing to the stress function $F$ that satisfy Eq. (7) Michell (1899), we retain only those that do not imply any dependence on $\varphi$ in the physical solution. We formulate the stress function separately for the inner and outer regions,
@@
II.3 Thickness of the shell for mechanical concealment
We now turn to the central point. The shell around the hole shall mechanically conceal and mask the hole in a way that its presence is not noted from outside. This concealment shall be realized for given materials, that is, we may not modify the material parameters $\mu^{\{\mathrm{i,o}\}}$ and $\nu^{\{\mathrm{i,o}\}}$. For that purpose, we must choose the outer radius $b$ of the shell in a way that the mechanical solution outside the shelled hole appears in the same way as if the hole were not present at all.
@@ II.3 Thickness of the shell for mechanical concealment
If the Poisson ratio of the shell is smaller compared to the surrounding material, $\nu^{\mathrm{i}}<\nu^{\mathrm{o}}$, the necessary outer radius and thus thickness of the shell diverges even with shells stiffer than the outside elastic material, $\mu^{\mathrm{i}}>\mu^{\mathrm{o}}$, if the shells are not overly stiff, see the red curves in Fig. 2. In such cases, for given materials, mechanical shielding will be challenging, and a stiffer choice of shell material may still be advisable. For example, the combination $\nu^{\mathrm{i}}=0.3$ and $\nu^{\mathrm{o}}=0.49$, see Fig. 2(a), depicts a compressible shell such as steel Greaves et al. (2011); Mott and Roland (2009) within a nearly incompressible background like a rubbery elastomer Greaves et al. (2011); Mott and Roland (2009). However, given that steels are generally several orders of magnitude stiffer than elastomers, this does not represent an actual constraint for mechanical concealment. It would still be possible already using very thin shells.
@@
Thus, where possible, selecting most beneficial materials provides advantages. Leveraging the existence of various unconventional materials, one may even opt for materials of near-zero or negative Poisson ratios Soman et al. (2012); Jiang and Park (2014); Yu, Yan, and Ruzsinszky (2017); Liu and Zhang (2018); Lee et al. (2019); Gaal et al. (2020); Gao et al. (2024) to achieve advanced mechanical shielding. However, once a choice is made or the materials are determined by other factors, then still for many practical material combinations the strategy of simply adjusting the thickness of the shell provides an effective, realizable way for mechanical concealment.
III Molecular dynamics simulations of mechanical cloaking
@@
To proceed a step beyond continuum considerations, we seek to investigate whether our analytical solution for mechanical concealment of a hole carries over to atomistic levels. If consistency between continuum and atomistic solutions is identified, it will indicate a robust cloaking strategy for large-scale applications. Thus, a scope of microscopic realization and adjustment of atomistic mechanical properties is conceivable. This correlation would allow us to connect continuum elasticity with microscopic structural stability of matter in the context of cloaking. To this end, we analyze a microscopic model solid using molecular dynamics (MD) simulations. For given combinations of $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$ and $b/a$. we test whether mechanical concealment can be achieved.
@@
In fact, atomistic investigations of mechanically concealed holes inside a solid are rare and challenging. Specifically, stabilizing solids with voids or cavities on the atomic scale often needs many-body terms in the interatomic interaction potentials. They come at a cost of losing simple, transparent models. Often, the descriptions invoke anisotropy due to angular dependence of the potentials. Our scope is to circumvent this issue and work with a description as simple as possible. Therefore, we use a two-dimensional, truncated and shifted Lennard-Jones (LJ) potential and adjust its length and energy scale, that is, the depth of the energy well, see Appendix A for details, even if resulting energy wells are significantly deeper than for standard rare-gas solids. They imply very high mechanical moduli of the order of TPa. As an advantage, we can work with a very simple and accessible potential and thus mimic a “toy model” for a strongly cohesive solid. As a further benefit, the LJ potential in two dimensions leads to stable hexagonal solids. The linearly elastic properties of solids of hexagonal symmetry are the same as for isotropic ones from a continuum perspective Landau and Lifshitz (1999). Therefore, our atomistic approach is in line with the continuum considerations above for isotropic elastic materials.
III.1 Simulation details and protocols
Molecular dynamics (MD) simulations are carried out in two dimensions using 46200 atoms interacting via a truncated and force-shifted Lennard-Jones (LJ) potential, see Fig. 10. It is defined as (Toxvaerd and Dyre, 2011)
$\begin{split}V_{SF}(r)=\begin{cases}V_{LJ}(r)-V_{LJ}(r_{c})\\ \quad-(r-r_{c})V_{LJ}^{\prime}(r_{c}),&r<r_{c},\\[6.0pt] 0,&r\geq r_{c},\end{cases}\end{split}$ (35)
with
$V_{LJ}(r)=4\epsilon\left[\left(\frac{\sigma}{r}\right)^{12}-\left(\frac{\sigma}{r}\right)^{6}\right],$ (36)
where $V_{SF}(r)$ and $V_{LJ}(r)$ are the force-shifted and the standard Lennard-Jones interatomic potentials, respectively. Here, $r$ denotes the center-to-center distance between two atoms. $r_{c}$ denotes the cutoff distance of the potential (=2.5=2.5 Å). As described in the main text, to mimic a “toy model” of a strongly cohesive two-dimensional background solid, $\epsilon=\epsilon^{\mathrm{o}}$ = 1.0 eV and $\sigma$ = 1.0 Å have been chosen as energy and length scale parameters respectively. Within the shell surrounding the hole, we set $\epsilon=\epsilon^{\mathrm{i}}$. The Lorentz-Berthelot mixing rule ($\epsilon=\sqrt{\epsilon^{\mathrm{i}}\epsilon^{\mathrm{o}}}$) is used to define the energy parameter of the potential between an atom belonging to the shell and an atom being part of the surrounded elastic solid. $\sigma$ = 1.0 Å is considered for all atoms in the solid.
[…]
For hexagonal solids with holes, mechanical concealment is achieved by introducing a concentric circular ring-like region (shell) of variable thickness and variable stiffness around the hole. A Lennard-Jones potential to realize our model solid is particularly useful in this context. In this case, we can directly relate the stiffness in terms of the shear modulus $\mu$ to the LJ energy parameter $\epsilon$ via $\mu$ $\approx$ ${\epsilon}/{r_{0}^{2}}$, with $r_{0}$ denoting the lattice constant (see Section III.2 and Fig. 12 in Appendix for details). Further, these pristine as well as hexagonal solids with holes (with and without shielding shell around the hole) are subjected to isotropic compression tests.
Holes were carved out mostly with a radius $a$ of 20 Å. However, several other hole radii or ratios of hole radius $a$ to the system size $L$ ($0.009<a/L<0.09$) are considered for a particular given ratio of the mechanical moduli $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$ in order to study the effect of $a/L$ on the compressive response of the systems. To this end, computational isotropic compression tests are performed on the equilibrated samples via unbiased area contraction using a constant engineering strain rate of $10^{-4}$ per timestep. We run these simulations for $5\times 10^{6}$ MD steps (corresponding to 0.5 ns) within NVT ensemble, out of which only the initial small-strain, linear pressure-areal strain regimes (up to 0.3 $\%$ areal strain) are used to compute the bulk modulus $K$ in two dimensions via
Figure 3: (a) Illustration of a representative isotropic compression test in an MD simulation of a planar hexagonal solid with a shielded hole ($b/a=2.018$, $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}=2$). Isotropic compression can be realized by imposing an isotropic stress $\bm{\sigma}^{\infty}=\bm{\sigma}(r\rightarrow\infty)=-P\mathbf{I}$, see the main text. (b) Isotropic compression is confirmed by plotting the components of the imposed stress $\sigma^{\infty}_{xx}$ and $\sigma^{\infty}_{yy}$ against each other. The linear fit confirms an approximate unit slope.
$K=-A_{0}\left(\frac{\partial P}{\partial A}\right)_{T}=-\left(\frac{\partial P}{\partial\epsilon_{A}}\right)_{T},$ (37)
where we defined the areal strain $\epsilon_{A}={\Delta A}/{A_{0}}$. Here, $A_{0}$ denotes the equilibrated area and $\Delta A$ represents the change in area from the reference area $A_{0}$ during isotropic compression to the new area $A$.
In Eq. (37), the negative sign signifies decreasing area under compression (increasing pressure). The setting for a typical isotropic compression test is illustrated in Fig. 3(a) for an equilibrated system with a shielded hole. For the same sample, isotropic compression can be confirmed from the unit slope between the $xx$- and $yy$-components of the imposed stress, see Fig. 3(b).
Mechanical concealment of holes is finally achieved using the following protocol. First, we choose a given stiffness of the shells surrounding the holes using a specific ratio of shear moduli $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$. Then, we tune the thickness of the shell by suitably varying the ratio between the outer radii of shell and hole $b/a$. The concealment ratio is the desired $b/a$ that gives identical bulk modulus ($K$) of the solid with hole to that of the pristine solid without any hole (within a relative error of $\leq 0.07\%$). Throughout our MD simulations we consider $\nu^{\mathrm{i}}$ $\approx$ $\nu^{\mathrm{o}}$ = 0.34 (see Section III.3 for details). All molecular dynamics simulations described in this article are performed using LAMMPS (Thompson et al., 2022).
III.2 Controlling the shear modulus of the shell
According to Ref. (Fisher, Halperin, and Morf, 1979), the elastic moduli of an isotropic, homogeneous lattice with two Lamé coefficients $\mu$ and $\lambda$ under short-range potentials can be obtained from the long-wavelength dispersion relations via
$m\omega_{T}^{2}=\mu(qr_{0})^{2},$ (38)
$m\omega_{L}^{2}=(\lambda+2\mu)(qr_{0})^{2}.$ (39)
Here, $\omega_{L,T}$ are the longitudinal and transverse phonon frequencies, $q$ denotes the wave number, $m$ represents the effective mass, and $r_{0}$ is the equilibrium lattice parameter.
As shown in Ref. (Zhang and Nelson, 2021), for particles in a hexagonal lattice interacting via LJ potentials, long-wavelength dispersion curves yield
$m\omega_{T}^{2}=\epsilon\frac{27}{r_{0}^{2}}(qr_{0})^{2},$ (40)
$m\omega_{L}^{2}=\epsilon\frac{27\times 3}{r_{0}^{2}}(qr_{0})^{2}.$ (41)
Therefore, comparing Eqs. (38)–(41), we find
$\lambda=\mu=27\frac{\epsilon}{r_{0}^{2}}.$ (42)
Here, $\epsilon$ is the LJ energy parameter. $r_{0}$ is fixed to 1.12 Å in our simulations, see Fig. 12 in Appendix. Thus, the relation between the ratios of shear moduli and the LJ energy parameters for shell and surrounding solid becomes $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}=\epsilon^{\mathrm{i}}/\epsilon^{\mathrm{o}}$.
III.3 Poisson ratio evaluation
In a hexagonal lattice composed of atoms interacting via an LJ potential, we noted for the Lamé coefficients $\lambda=\mu$. Therefore, the Poisson ratio becomes $\nu=\lambda/(2\mu+\lambda)=1/3$. Indeed, from our MD simulations, we extract $\nu=0.34$ from the slope of the linear variation between transverse strain ($\epsilon_{yy}$) and applied strain ($\epsilon_{xx}$) via uniaxial tensile test of the equilibrated pristine hexagonal solid, see Fig. 4(a). As shown in Fig. 4(b), introducing a stiff shell surrounding a hole does not significantly alter $\nu$ in our case, here to a value of approximately 0.35. Throughout the evaluations of our MD simulations, we thus assume $\nu^{\mathrm{i}}$ $\approx$ $\nu^{\mathrm{o}}$ = 0.34.
Figure 4: Poisson ratios ($\nu$), obtained from the slope of the linear variation between an applied strain ($\epsilon_{xx}$) and a resulting transverse strain ($\epsilon_{yy}$) via performing uniaxial tensile tests of the equilibrated samples. Results are shown for (a) a pristine planar solid and (b) a planar solid with a hole enclosed by a stiffer shell of ratios of outer radii $b/a=1.173$ and of shear moduli $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}=10$.
III.4 MD results: shell thickness for concealment
Pristine hexagonal solids, hexagonal solids with a hole but no shell ($b/a=1$), and hexagonal solids containing shelled holes ($b/a>1$) were prepared. We here use the bulk modulus to compare their mechanical behavior and test the effective mechanical shielding of the hole from isotropic compression.